OpenStudy (anonymous):

2(x-6)^2=24 solve the equation. round the solutions to the nearest hundredth, if necessary.

OpenStudy (anonymous):

\[2(x-6)^2=24\] divide both sides by 2:\[(x-6)^2=12\] take square root:\[x-6=\sqrt{12}\] or \[x-6 = -\sqrt{12}\] therefore:\[x = 6+\sqrt{12}\] or \[x=6-\sqrt{12}\]